# Decoding the AI Pen: Techniques and Challenges in Detecting AI-Generated Text

### v1

@@ 5. Possibility of Detection Through a Theoretical Lens
where $TV(\mathcal{M},\mathcal{H})$ is the total variation distance between machine and human-written texts. This formula indicates that when human and machine generated texts are very similar i.e., $TV(\mathcal{M},\mathcal{H})$ is very small, even the most effective detector may exhibit only marginal improvement over a random classifier.
However, another interesting study by Chakraborty et al. (Chakraborty et al., 2023) suggests that as long as the distributions of human-generated and AI-generated texts are not identical (which is typically the case), it remains feasible to detect AI-generated texts. This detection becomes possible when we gather sufficient samples from each distribution. Interestingly, Chakraborty et al. demonstrate that the AUROC curve, as proposed by Sadasivan et al., might be overly conservative for practical detection scenarios. They introduce a hidden possibility by replacing $TV(\mathcal{M}^{,}\mathcal{H})$ with $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ in AUROC equation, where $m^{\bigotimes n}:=m\bigotimes m\bigotimes\dots\bigotimes m$ (n times) denotes the product distribution over sample set $\mathcal{S}:=\{s_{i}\},i\in\{1,\dots n\}$, as does $h^{\bigotimes n}$. Since $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ is an increasing sequence, it eventually converges to 11 as the number of samples for each distribution increases. If this happens, the total variation distance approaches 1 quickly, and hence increasing the AUROC.
@@
The authors introduce two key assumptions in their work. First, they propose the concept of a “Quality Oracle”, which grants the attacker access to an oracle capable of evaluating the quality of outputs. This oracle assists the attacker in assessing the quality of modified responses. Second, they introduce the “Perturbation Oracle”, which allows the attacker to modify an output while maintaining a nontrivial probability of preserving quality. The perturbation oracle essentially induces an efficiently mixing random walk on high-quality outputs. Their investigation culminates in a compelling result: For every public or secret-key watermarking setting that satisfying these assumptions, there exists an efficient attacker. “Given a prompt $p$ and a watermarked output $y$, this attacker can leverage the quality and perturbation oracles to obtain an output ${y}^{{}^{\prime}}$ with a probability very close to 11. The attacker’s goal is to find an output ${y}^{{}^{\prime}}$ such that (1) ${y}^{{}^{\prime}}$ is not watermarked with high probability and (2) ${Q(p,{y}^{{}^{\prime}})}\geq{Q(p,y)}$” (Zhang et al., 2023). In simpler terms, watermarking without causing significant quality degradation is impossible. Thus, the authors propose alternative approaches to safeguard the intellectual property of model developers.
@@ 6. New Opportunities and Future Research
This paper offers a comprehensive overview of the latest advancements and recommended practices in detecting AI-generated text. However, existing methods often face limitations and are susceptible to malicious attacks. Thus, it is essential to continue exploring both theoretical and practical aspects of this realm. In light of this, we propose opportunities for advancing the field of responsible AI, vulnerability assessment, and risk mitigation studies related to LLMs. Some of these opportunities include:
Enhancing the diversity and representativeness of datasets is crucial for training and evaluating AI-text detection models. The existing datasets may not comprise all the potential types and sources of AI-generated content.
Further investigating interpretable features to discern the subtle differences between human-written and AI-generated text considering the limited availability of such features.
@@
A comprehensive understanding of the fundamental feasibility and boundaries as there is often an absence of a thorough grasp of the fundamental feasibility and limitations within current SOTA methods. Therefore, there is a need for deeper exploration and investigation into the theoretical aspects of this task to facilitate the creation of more resilient and efficient techniques.

### v2

@@ 4.1. Impossibility via AUROC Upper-bound
where $TV(\mathcal{M},\mathcal{H})$ is the total variation distance between machine and human-written texts distributions. This formula indicates that when human and machine generated texts are very similar i.e., $TV(\mathcal{M},\mathcal{H})$ is very small, even the most effective detector may exhibit only marginal improvement over a random classifier. Figure 3 illustrates how the above bound grows as a function of the total variation distance.
Figure 3. Comparing AUROC of the optimal detector to a random classifier demonstrates that as the TV distance between AI and human text distributions reduces, the AUROC of the optimal detector also decreases accordingly.
4.2. Possibility via Sample Abundance
Another interesting study by Chakraborty et al. (Chakraborty et al., 2023) suggests that:
As long as the distributions of human-generated and AI-generated texts are not identical (which is typically the case), it remains feasible to detect AI-generated texts. This detection becomes possible when we gather sufficient samples from each distribution.
Interestingly, Chakraborty et al. demonstrate that the AUROC curve, as proposed by Sadasivan et al., might be overly conservative for practical detection scenarios. Specifically, they introduce a hidden possibility by replacing $TV(\mathcal{M},\mathcal{H})$ with $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ in AUROC equation, where $m^{\bigotimes n}:=m\bigotimes m\bigotimes\dots\bigotimes m$ (n times) denotes the product distribution over sample set $\mathcal{S}:=\{s_{i}\},i\in\{1,\dots n\}$, as does $h^{\bigotimes n}$. For any detector $\mathcal{D}$, with a given collection of i.i.d. samples either from human or AI, the AUROC curve is:
$AUROC(\mathcal{D})\leq\frac{1}{2}+TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})-\frac{TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})^{2}}{2}$
where $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})=1-exp(-nl_{c}(m,h)+o(n)$ and $l_{c}(m,h)$ is the Chernoff information. Since $TV(\mathcal{M}^{\bigotimes n},\mathcal{H}^{\bigotimes n})$ is an increasing sequence, it eventually converges to $1$ as the number of samples for each distribution increases. Therefore, the upper bound of AUROC increases exponentially with respect to the number of samples. It is clear that if this happens, the total variation distance approaches $1$ quickly, and hence increasing the AUROC. Interested readers are encouraged to refer to (Chakraborty et al., 2023) for further details.
4.3. Impossibility of Strong Watermarking
@@
Zhang et al. put forth two fundamental assumptions:
“Quality Oracle”:
grants the attacker access to an oracle capable of evaluating the quality of outputs. This oracle assists the attacker in assessing the quality of modified responses.
“Perturbation Oracle”:
allows the attacker to modify an output while maintaining a nontrivial probability of preserving quality.
The perturbation oracle essentially induces an efficiently mixing random walk on high-quality outputs. Their investigation culminates in a compelling result:
Given a prompt $p$ and a watermarked output $y$, for every public or secret-key watermarking setting that satisfying these assumptions, there exists an efficient attacker that can leverage the quality and perturbation oracles to obtain an output ${y}^{{}^{\prime}}$ with a probability very close to $1$. The attacker’s goal is to: find an output ${y}^{{}^{\prime}}$ s.t. (1) ${y}^{{}^{\prime}}$ is not watermarked with high probability and, (2) ${Q(p,{y}^{{}^{\prime}})}\geq{Q(p,y)}$” (Zhang et al., 2023).
In simpler terms, watermarking without causing significant quality degradation is impossible. Thus, Zhang et al. propose alternative approaches to safeguard the intellectual property of developers.
@@ 5. New Opportunities and Future Research
This paper offers a comprehensive overview of the latest advancements and recommended practices in detecting AI-generated text. However, existing methods often face limitations and are susceptible to malicious attacks. Thus, it is essential to continue exploring both theoretical and practical aspects of this realm. In light of this, we propose opportunities for advancing the field of responsible AI, vulnerability assessment of AI-generated text detection studies. Some of these opportunities include:
  • Curating diverse and representative datasets is crucial for training and evaluating detection models. The existing datasets may not comprise all the potential types and sources of AI text. Specially, it is essential to curate a dataset containing outputs from various generative models.
  • Further investigating interpretable features can help discern the subtle differences between human-written and AI-generated text, especially given the limited availability of such features. Additionally, investigating their vulnerability to adversarial attacks is crucial.
@@
  • A comprehensive multi-aspect evaluation of detection techniques against adversarial attacks as literature lacks such study to investigate relative vulnerabilities of detection methods. Such an analysis should consider various perspectives, including efficacy across different models and resilience against adversarial attacks. This study can aid in identifying techniques suitable for specific scenarios.
  • Developing hybrid detection strategies that combine features and techniques to enhance robustness and adaptability. For example, one can create a hybrid approach by integrating watermarking and feature-based techniques.
  • A comprehensive understanding of the fundamental feasibility and boundaries as there is often an absence of a thorough grasp of the fundamental feasibility and limitations within current state-of-the-art methods. Therefore, there is a need for deeper exploration and investigation into the theoretical aspects of this task to facilitate the creation of more resilient and efficient techniques while also uncovering novel avenues for research.
@@ 6. Conclusions
{tblr}
[!t] colspec = @p3cmp2.25cmp5.5cmp5.5cm, row1 = bg=blue!20, row2 = bg=blue!3, row3 = bg=blue!3, row4 = bg=blue!3, row5 = bg=blue!3, row6 = bg=blue!3, Related PapersMethodMain IdeaVulnerabilities
(Solaiman et al., 2019; Bakhtin et al., 2019; Antoun et al., 2023; Zhan et al., 2023; Li et al., 2023) Supervised
detectionTo fine-tune a model on sets of AI and human generated texts. Training on commonly used datasets, makes it vulnerable to most attacks including paraphrasing.
 (Su et al., 2023; Zer, 2023; Mitchell et al., 2023; Gehrmann et al., 2019; Wang et al., 2023a; Guo and Yu, 2023) Zero-shot
detectionTo use a pre-trained language model in zero-shot settings. Reduces the risk of data poisoning attacks and eliminates data and resource over- heads, but it is still susceptible to other adversarial attacks like spoofing and paraphrasing.
 (Krishna et al., 2023; Sadasivan et al., 2024; Wolff, 2020; Liang et al., 2023) Retrieval-based
detection To apply methods of information retrieval to match a given text with a collection of texts generated by LLMs and finding similarities in meaning. It is impractical because it requires a large and updated collection of texts, which is computationally expensive, or may be unavailable for all domains, tasks or models. It is also vulnerable to paraphrasing and spoofing attacks.
(Kirchenbauer et al., 2023a; Yang et al., 2023a; Kirchenbauer et al., 2023b; Sadasivan et al., 2024) WatermarkingTo use a model signature in the produced text outputs to stamp particular pattern. The most trustworthy strategy, but is shown to be fundamentally impossible for generative models. It is susceptible to attacks such as rewording and spoofing.
(Yu et al., 2023; Yang et al., 2023b; Mitchell et al., 2023; Su et al., 2023; Zhang et al., 2023) Feature-based
detectionTo identify and classify based on extracted discriminating features.Susceptible to adversarial attacks such as paraphrasing.
%bibliographystyleACM-Reference-Format.bst